# Existence of Minimal Surfaces in Infinite Volume Hyperbolic $3$-manifolds

### v1

@@ Abstract.
Among other results, for doubly degenerate manifolds with bounded geometry, we prove a dichotomy: either every such manifold contains a closed minimal surface or there exists such a manifold admitting a foliation by closed minimal surfaces. We also construct the first examples of Schottky manifolds with closed minimal surfaces and demonstrate the existence of Schottky manifolds containing infinitely many closed minimal surfaces. Lastly, for hyperbolic 3-manifolds with rank-1 cusps, we show that a broad class of these manifolds must contain a finite-area, embedded, complete minimal surface.
@@ 1. Introduction and Statement of Main Results
The existence of minimal surfaces in Riemannian $3$-manifolds is one of the classical problems in geometric analysis. Significant progress has been made in understanding the existence of closed embedded minimal surfaces in compact 3-manifolds, particularly in the positive curvature case. Using the Min-Max approach, Pitts and Simon-Smith established that every closed Riemannian 3-manifold contains a smooth, embedded, closed minimal surface [Pit14], [Smi83], [CDL03]. Through the methods of geometric measure theory, Federer [Fed14] demonstrated that for any closed Riemannian 3-manifold, an area-minimizing surface exists in each homology class. Meeks, Simon, and Yau [MSY82] further showed that every isotopy class contains a minimal surface in such manifolds. More recently, Song [Son23b] resolved Yau’s Conjecture, proving that any closed 3-manifold contains infinitely many minimal surfaces, building on techniques developed by Marques and Neves [MN17].
@@
Next, we study the missing cases above. For case (1) - product manifolds, we prove that a strong dichotomy holds when $\mathcal{M}$ has bounded geometry, or a uniform lower bound for its injectivity radius at every point:
Theorem 1.2 (Dichotomy for Doubly Degenerate Manifolds).
@@
This theorem establishes a connection between the existence theory of minimal surfaces and the longstanding Anderson-Thurston conjecture, on the non-existence of minimal foliations of hyperbolic 3-manifolds (see Question 7.1 below.) Building on the ideas of the proof of this result, we show that the growth rate of disjoint stable minimal surfaces is controlled by the volume of the convex core, conditional on a mild strengthening of the Anderson-Thurston conjecture. We are able to prove an unconditional statement for certain quasi-Fuchsian manifolds with short geodesics (4.6), and we provide an explicit construction of such manifolds.
@@
For case (3) — manifolds with rank-1 cusps — we use shrinkwrapping surfaces in the ends [CG06] to show the existence of a complete, finite-area, embedded minimal surface for a broad class of such manifolds. While our results are more general (see Section 6.2), here we highlight an important class of manifolds that this result covers.
@@
Let $\mathcal{M}$ be an infinite volume complete hyperbolic $3$-manifold with rank-1 cusps. Assume $\mathcal{M}$ is topologically not a handlebody, nor geometrically infinite product manifold. If one of the ends is geometrically finite, then $\mathcal{M}$ contains a finite area, complete, embedded, minimal surface $\mathcal{S}$.
@@
In Section 2, we overview the related results and give a basic background on the geometry and topology of hyperbolic $3$-manifolds. In Section 3, we prove the thickness at infinity result for non-cusped hyperbolic $3$-manifolds, and prove 1.1. In Section 4, we first establish our dichotomy result for doubly degenerate manifolds (1.2). Next, we prove our dichotomy result for quasi-Fuchsian manifolds (1.3). In Section 5, we give the construction of the first examples of Schottky manifolds with closed minimal surfaces (1.4). In Section 6, we prove the existence of finite area, complete, embedded, minimal surfaces in hyperbolic $3$-manifolds with rank-1 cusps (1.5 and 1.3). In Section 7, we give our final remarks and discuss the remaining open questions.
@@ 2.1. Finite Volume Hyperbolic $3$-Manifolds
There exist two types of finite volume hyperbolic $3$-manifolds, classified by whether or not they are compact: compact (closed) and non-compact (cusped). The existence of closed embedded minimal surfaces in closed hyperbolic $3$-manifolds was established decades ago via Almgren-Pitts min-max theory [Pit14]. Note that finite volume cusped hyperbolic $3$-manifolds only have rank-2 cusps [BP92]. This implies that the manifold is homeomorphic to the interior of a compact $3$-manifold where every boundary component is a 2-torus. Recently, the existence of closed minimal surfaces has been shown in this non-compact case by several authors [CHMR17, HW17, Rub87]. Furthermore, Chambers-Liokumovich obtained a very general existence result for any finite volume noncompact Riemannian $3$-manifold [CL20] by generalizing Almgren-Pitts min-max techniques. Recently, Song showed that finite volume $3$-manifolds contain infinitely many closed embedded minimal surfaces [Son23a].
@@ 3. Dichotomy for Non-Cusped Hyperbolic $3$-manifolds
We note that the above results and Song’s dichotomy theorem shed no light on the number of stable minimal surfaces in a manifold $\mathcal{M}$. A manifold can have a stable minimal surface without closed saddle point minimal surfaces. For example, let $\mathcal{M}\simeq\mathcal{S}\times\mathbf{R}$ be a Fuchsian manifold, then $\mathcal{M}$ contains a unique minimal surface, which is the totally geodesic surface $\mathcal{S}\times\{0\}$. More generally, every almost Fuchsian manifold [Uhl83] only contains exactly one closed minimal surface, and there are also examples that are not almost Fuchsian [HLar].
@@
We note that our thickness at infinity result does not generalize to hyperbolic 3-manifolds with cusps. For this case, a straightforward example is a Fuchsian manifold ×$\mathcal{F}\times\mathbf{R}$ where $\mathcal{F}$ is a punctured surface. We provide further examples of cusped, quasi-Fuchsian hyperbolic $3$-manifolds containing finite-area, non-compact, complete, embedded minimal surfaces in 6.14.
4. Product Manifolds
In this section, we first present our dichotomy result for doubly degenerate  hyperbolic $3$-manifolds with bounded geometry. These manifolds can be seen as limits of quasi-Fuchsian manifolds. As an application, we obtain results on counting minimal surfaces in these manifolds. In particular, we derive a linear growth estimate for the number of disjoint embedded minimal surfaces in quasi-Fuchsian manifolds with short geodesics (Section 4.2), and we provide an explicit construction of these manifolds (Section 4.3).
4.1. Dichotomy for Doubly Degenerate Manifolds
@@
The condition locally area-minimizing is essential here. Using the fact that the function $f(t)=t^{3}$ vanishes to second order at 00, it is possible to construct Riemannian manifolds $M$ for which the conclusion of the lemma can fail if $S$ is only assumed to be stable. Furthermore, the work in [GLVP21] shows that it can also fail for $M$ hyperbolic and satisfying the assumptions of the lemma.
Now, we are ready to state our dichotomy result.
@@
For each $\sigma\in\text{Teich}(S)$, denote the corresponding harmonic maps $S\rightarrow\mathcal{M}$ and $S\rightarrow\mathcal{M}_{\infty}$ by $F_{\sigma}$ and $F_{\sigma}^{\infty}$. Then provided that the energy $E(F_{\sigma})$ is bounded from above by some constant $C$, it will be the case that $E(F_{G_{n}(\sigma)}^{\infty})$ is bounded from above by $C+\delta(n)$ for $\delta(n)$ tending to zero as $n$ tends to infinity, by the fact that the $\Phi_{n}$ are smoothly converging to isometries on compact sets. Here we also use the fact that the image of any of the harmonic maps that we are considering has diameter bounded above by a constant depending only on $\mathcal{M}$. In [Min93][Lemma 3.1], it is shown that each harmonic map has a bound on the diameter of its image depending only on the injectivity radius lower bound for $\mathcal{M}$ and the topological type of $\mathcal{M}$.
@@
We now apply 4.1 to $\mathcal{M}_{\infty}$. We just need to check that if we are in the second alternative of 4.1, then $\mathcal{M}$ as in the previous subsection has infinitely many minimal surfaces. In this case there are surfaces $S_{\epsilon}$ and $S_{-\epsilon}$ that bound a product region homeomorphic to $S\times(-\epsilon,\epsilon)$ in $\mathcal{M}$ containing $S$, and whose mean-curvature vectors point towards $S$ at every point. In this case, for large $n$ we have that $\Phi_{n}(S_{\pm\epsilon})$ bound a product region in $\mathcal{M}$ with mean-convex boundary. Each such product region contains a locally area-minimizing minimal surface by [MSY82], which gives infinitely many minimal surfaces in $\mathcal{M}$. Using min-max we can then produce infinitely many “saddle” minimal surfaces as well.
@@
Our dichotomy result when combined with Gromov’s dichotomy [Gro14], gives an interesting relation. In particular, Gromov’s dichotomy states that a complete noncompact manifold either contains a finite area minimal surface or there exists a strictly mean convex foliation of every compact subdomain. Assuming no rank-1 cusps in $\mathcal{M}$, 4.3 implies that if there is a doubly degenerate manifold $\mathcal{M}$ with bounded geometry which admits a strictly convex mean-convex foliation (hence no closed minimal surface by Gromov’s dichotomy), then there must be another such manifold $\mathcal{M}$ with a minimal foliation.
@@ 4.3. Quasi-Fuchsian Manifolds with Short Geodesics
Assume that $M=\Sigma\times\mathbb{R}$ is a quasi-Fuchsian manifold with more than one closed local area minimizer. Without loss of generality, we can assume these local area minimizers $\Sigma,\Sigma^{\prime}$ are disjoint. Indeed, if they intersect, we can minimize area in each unbounded component of their complement to produce a disjoint pair. By Corollary 3.5 it is sufficient to prove that there exists a saddle point minimal surface. Let $\Omega$ be the compact region bounded by $\Sigma,\Sigma^{\prime}$ and take $\Sigma=\Sigma_{1},\ldots,\Sigma_{n}=\Sigma^{\prime}$ a maximal collection of stable minimal surfaces isotopic to $\Sigma\times\{0\}$ in the region $\Omega$, so that each consecutive pair bounds a cylinder disjoint from any of the other surfaces. Without loss of generality assume that all surfaces in this maximal collection are not saddle. Then at least one consecutive pair satisfies the conditions of [GLVP21, Proposition 7.6], from which the existence of a saddle point minimal surface follows. ∎
@@
There are examples of either cases in Corollary 4.8. As mentioned before, almost Fuchsian manifolds ([Uhl83, HW13]) satisfy condition (1). For examples satisfying (2) one can take the manifolds from [HW19]. Alternatively, it is known that there exist quasi-Fuchsian manifolds with at least one degenerate stable minimal surface homotopic to the fiber (see for instance [GLVP21, Theorem 1.6] where that is shown to occur for outermost minimal surfaces). By work of Uhlenbeck ([Uhl83, Theorem 4.4, Corollary 5.2]) we can deform such quasi-Fuchsian manifold to quasi-Fuchsian manifold with a strictly unstable minimal surface homotopic to the fiber. Such quasi-Fuchsian manifolds satisfy (2) by Corollary 3.5.
@@ 5. Schottky Manifolds
By using for instance the examples from 5.1, take any bounded geometry handlebody with two disjoint local area minimizers isotopic to the boundary. Applying the reasoning of [GLVP21, Proposition 7.6] as in Corollary 4.8, by the local min-max theorem [KLS19, Theorem 10] we have a saddle point minimal surface in between. Hence the result follows once more by applying 3.5. ∎

### v2

@@ Abstract.
Among other results, for doubly degenerate manifolds with bounded geometry, we prove an alternative: either every such manifold contains a closed minimal surface or there exists such a manifold admitting a foliation by closed minimal surfaces. We also construct the first examples of Schottky manifolds with closed minimal surfaces and demonstrate the existence of Schottky manifolds containing infinitely many closed minimal surfaces. Lastly, for hyperbolic 3-manifolds with rank-1 cusps, we show that a broad class of these manifolds must contain a finite-area, embedded, complete minimal surface.
@@ 1. Introduction and Statement of Main Results
The existence of minimal surfaces in Riemannian $3$-manifolds is one of the classical problems in geometric analysis. Significant progress has been made in understanding the existence of closed embedded minimal surfaces in compact 3-manifolds, particularly in the positive curvature case. Using the Min-Max approach, Pitts and Simon-Smith established that every closed Riemannian 3-manifold contains a smooth, embedded, closed minimal surface [Pit14], [Smi83], [CDL03]. By the methods of geometric measure theory, Federer [Fed14] demonstrated that for any closed Riemannian 3-manifold, an area-minimizing surface exists in each homology class. Meeks, Simon, and Yau [MSY82] further showed that every isotopy class contains a minimal surface in such manifolds. More recently, Song [Son23b] resolved Yau’s Conjecture, proving that any closed 3-manifold contains infinitely many minimal surfaces, building on techniques developed by Marques and Neves [MN17].
@@
Next, we study the missing cases above. For case (1) - product manifolds, we prove that an alternative holds when $\mathcal{M}$ has bounded geometry, or a uniform lower bound for its injectivity radius at every point:
Theorem 1.2 (Alternative for Doubly Degenerate Manifolds).
@@
This theorem establishes a connection between the existence theory of minimal surfaces and the longstanding Anderson-Thurston conjecture, on the non-existence of minimal foliations of hyperbolic 3-manifolds (see Question 7.1 below). We note that the two possibilities in the alternative are not mutually exclusive.
Building on the ideas of the proof of this result, we show that the growth rate of disjoint stable minimal surfaces is controlled by the volume of the convex core, conditional on a mild strengthening of the Anderson-Thurston conjecture. We are able to prove an unconditional statement for certain quasi-Fuchsian manifolds with short geodesics (4.6), and we provide an explicit construction of such manifolds.
@@
Finally, we show the existence of a complete, finite-area, embedded minimal surface for a broad class of hyperbolic 3-manifolds with rank-1 cusps. In contrast to the first author’s previous work on the existence problem for minimal surfaces using shinkwrapping surfaces [Cos21], our proof uses a doubling trick that depends on an analysis of the behavior of the convex core near the rank one cusps.
@@
Let $\mathcal{M}$ be an infinite volume complete hyperbolic $3$-manifold with rank-1 cusps homeomorphic to the interior of a compact manifold with boundary. Assume $\mathcal{M}$ is topologically neither a handlebody nor a geometrically infinite product manifold. If one of the ends of $\mathcal{M}$ is geometrically finite, then $\mathcal{M}$ contains a finite area, complete, embedded, minimal surface $\mathcal{S}$.
Remark 1.6.
A compact, oriented, irreducible, atoroidal 3-manifold with non-empty boundary carries many infinite volume hyperbolic metrics that have a geometrically finite end with a rank-1 cusp, see e.g. [CCHS03].
@@
In Section 2, we overview the related results and give a basic background on the geometry and topology of hyperbolic $3$-manifolds. In Section 3, we prove the thickness at infinity result for non-cusped hyperbolic $3$-manifolds, and prove 1.1. In Section 4, we first establish our alternative result for doubly degenerate manifolds (1.2). Next, we prove our dichotomy result for quasi-Fuchsian manifolds (1.3). In Section 5, we give the construction of the first examples of Schottky manifolds with closed minimal surfaces (1.4). In Section 6, we study hyperbolic $3$-manifolds with rank-1 cusps and prove 1.5. In Section 7, we give our final remarks and discuss the remaining open questions.
@@ 2.1. Finite Volume Hyperbolic $3$-Manifolds
There exist two types of finite volume hyperbolic $3$-manifolds, classified by whether or not they are compact: compact (closed) and non-compact (cusped). The existence of closed embedded minimal surfaces in closed hyperbolic $3$-manifolds was established decades ago via Almgren-Pitts min-max theory [Pit14]. Since finite volume cusped hyperbolic $3$-manifolds only have rank-2 cusps [BP92], this implies that the manifold is homeomorphic to the interior of a compact $3$-manifold where every boundary component is a 2-torus. Recently, the existence of closed minimal surfaces has been shown in this non-compact case by several authors [CHMR17, HW17, Rub87]. Furthermore, Chambers-Liokumovich obtained a very general existence result for any finite volume noncompact Riemannian $3$-manifold [CL20] by generalizing Almgren-Pitts min-max techniques. Recently, Song showed that finite volume $3$-manifolds contain infinitely many closed embedded minimal surfaces [Son23a].
@@ 3. Dichotomy for Non-Cusped Hyperbolic $3$-manifolds
We note that the above results and Song’s dichotomy theorem shed no light on the number of stable minimal surfaces in a manifold $\mathcal{M}$. A manifold can have a stable minimal surface without closed saddle point minimal surfaces. For example, let $\mathcal{M}\simeq\mathcal{S}\times\mathbf{R}$ be a Fuchsian manifold, then $\mathcal{M}$ contains a unique minimal surface, which is the totally geodesic surface $\mathcal{S}\times\{0\}$. More generally, every almost Fuchsian manifold [Uhl83] only contains exactly one closed minimal surface, and there are also examples that are not almost Fuchsian [HL25].
@@
We note that our thickness at infinity result does not generalize to hyperbolic 3-manifolds with cusps. For this case, a straightforward example is a Fuchsian manifold ×$\mathcal{F}\times\mathbf{R}$ where $\mathcal{F}$ is a punctured surface. We provide further examples of cusped, quasi-Fuchsian hyperbolic $3$-manifolds containing finite-area, non-compact, complete, embedded minimal surfaces in section six.
4. Product Hyperbolic $3$-Manifolds
In this section, we first present our alternative result for doubly degenerate  hyperbolic $3$-manifolds with bounded geometry. These manifolds can be seen as limits of quasi-Fuchsian manifolds. As an application, we obtain results on counting minimal surfaces in these manifolds. In particular, we derive a linear growth estimate for the number of disjoint embedded minimal surfaces in quasi-Fuchsian manifolds with short geodesics (Section 4.2), and we provide an explicit construction of these manifolds (Section 4.3).
4.1. Alternative for Doubly Degenerate Manifolds
@@
The condition locally area-minimizing is essential here. Using the fact that the function $f(t)=t^{3}$ vanishes to second order at 00, it is possible to construct Riemannian manifolds $M$ for which the conclusion of the lemma can fail if $S$ is only assumed to be stable. Furthermore, the work in [GLVP25] shows that it can also fail for $M$ hyperbolic and satisfying the assumptions of the lemma.
Now, we are ready to state the main results of this section.
@@
For each $\sigma\in\text{Teich}(S)$, denote the corresponding harmonic maps $S\rightarrow\mathcal{M}$ and $S\rightarrow\mathcal{M}_{\infty}$ by $F_{\sigma}$ and $F_{\sigma}^{\infty}$. Then provided that the energy $E(F_{\sigma})$ is bounded from above by some constant $C$, it will be the case that $E(F_{G_{n}(\sigma)}^{\infty})$ is bounded from above by $C+\delta(n)$ for $\delta(n)$ tending to zero as $n$ tends to infinity, by the fact that the $\Phi_{n}$ are smoothly converging to isometries on compact sets. Here we also use the fact that the image of any of the harmonic maps that we are considering has diameter bounded above by a constant depending only on $\mathcal{M}$: in [Min93][Lemma 3.1], it is shown that each harmonic map has a bound on the diameter of its image depending only on the injectivity radius lower bound for $\mathcal{M}$ and the topological type of $\mathcal{M}$.
@@
We now apply 4.1 to $\mathcal{M}_{\infty}$. We just need to check that if we are in the second alternative of 4.1, then $\mathcal{M}$ has infinitely many minimal surfaces. In this case there are surfaces $S_{\epsilon}$ and $S_{-\epsilon}$ that bound a product region homeomorphic to $S\times(-\epsilon,\epsilon)$ in $\mathcal{M}_{\infty}$ containing $S$, and whose mean-curvature vectors point towards $S$ at every point. In this case, for large $n$ we have that $\Phi_{n}(S_{\pm\epsilon})$ bound a product region in $\mathcal{M}$ with mean-convex boundary. Each such product region contains a locally area-minimizing minimal surface by [MSY82], which gives infinitely many minimal surfaces in $\mathcal{M}$ because the regions bounded by the $\Phi_{n}(S_{\pm\epsilon})$ will leave every compact set as $n$ tends to infinity. Using min-max we can then produce infinitely many “saddle” minimal surfaces as well.
@@
Our alternative result when combined with Gromov’s dichotomy [Gro14], gives an interesting relation. In particular, Gromov’s dichotomy states that a complete noncompact manifold either contains a finite area minimal surface or there exists a strictly mean convex foliation of every compact subdomain. Assuming no rank-1 cusps in $\mathcal{M}$, 4.3 implies that if there is a doubly degenerate manifold $\mathcal{M}$ with bounded geometry which admits a strictly convex mean-convex foliation (hence no closed minimal surface by Gromov’s dichotomy), then there must be another such manifold $\mathcal{M}$ with a minimal foliation.
@@ 4.3. Quasi-Fuchsian Manifolds with Short Geodesics
Assume that $M=\Sigma\times\mathbb{R}$ is a quasi-Fuchsian manifold with more than one closed local area minimizer. Without loss of generality, we can assume these local area minimizers $\Sigma,\Sigma^{\prime}$ are disjoint. Indeed, if they intersect, we can minimize area in each unbounded component of their complement to produce a disjoint pair. By Corollary 3.5 it is sufficient to prove that there exists a saddle point minimal surface. Let $\Omega$ be the compact region bounded by $\Sigma,\Sigma^{\prime}$ and take $\Sigma=\Sigma_{1},\ldots,\Sigma_{n}=\Sigma^{\prime}$ a maximal collection of stable minimal surfaces isotopic to $\Sigma\times\{0\}$ in the region $\Omega$, so that each consecutive pair bounds a cylinder disjoint from any of the other surfaces. Without loss of generality assume that all surfaces in this maximal collection are not saddle. Then at least one consecutive pair satisfies the conditions of [GLVP25, Proposition 7.6], from which the existence of a saddle point minimal surface follows. ∎
@@
There are examples of either cases in Corollary 4.8. As mentioned before, almost Fuchsian manifolds ([Uhl83, HW13]) satisfy condition (1). For examples satisfying (2) one can take the manifolds from [HW19]. Alternatively, it is known that there exist quasi-Fuchsian manifolds with at least one degenerate stable minimal surface homotopic to the fiber (see for instance [GLVP25, Theorem 1.6] where that is shown to occur for outermost minimal surfaces). By work of Uhlenbeck ([Uhl83, Theorem 4.4, Corollary 5.2]) we can deform such quasi-Fuchsian manifold to quasi-Fuchsian manifold with a strictly unstable minimal surface homotopic to the fiber. Such quasi-Fuchsian manifolds satisfy (2) by Corollary 3.5.
@@ 5. Schottky Manifolds
By using for instance the examples from 5.1, take any bounded geometry handlebody with two disjoint local area minimizers isotopic to the boundary. Applying the reasoning of [GLVP25, Proposition 7.6] as in Corollary 4.8, by the local min-max theorem [KLS19, Theorem 10] we have a saddle point minimal surface in between. Hence the result follows once more by applying 3.5. ∎